The following are the selling price, variable costs, and contribution margin for one unit of each of Banner Company’s three products: A, B, and C:

Product
A B C
  Selling price $ 110.00 $ 150.00 $ 170.00
  Variable costs:
    Direct materials 69.50 72.50 117.50
    Direct labour 6.00 20.00 8.00
    Variable manufacturing overhead 1.50 5.00 2.00
  Total variable cost 77.00 97.50 127.50
  Contribution margin $ 33.00 $ 52.50 $ 42.50
  Contribution margin ratio 30 % 35 % 25 %


Due to a strike in the plant of one of its competitors, demand for the company’s products far exceeds its capacity to produce. Management is trying to determine which product(s) to concentrate on next week in filling its backlog of orders. The direct labour rate is $4 per hour, and only 3,290 hours of labour time are available each week.


Required:
1.
Compute the amount of contribution margin that will be obtained per hour of labour time spent on each product. (Round your intermediate calculations to 1 decimal place. Round your answers to 2 decimal places.)

2. Which orders would you recommend that the company work on next week—the orders for product A, product B, or product C?

  • Product A

  • Product B

  • Product C


3. By paying overtime wages, more than 3,290 hours of direct labour time can be made available next week. Up to how much should the company be willing to pay per hour in overtime wages as long as there is unfilled demand for the three products? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

1.

Calculation of labour hour per unit

A = $6/4 = 1.5 hour

B = $20/4 = 5 hour

C = $8/4 = 2 hour

Contribution margin per hour = Contribution per unit / labor hour per unit

A = $33 / 1.5 hour = $22 per hour

B = $52.50 / 5 hour = $10.5 per hour

C = $42.50 / 2 hour = $10.625 per hour

2. Company should work on Product A first because it is providing highest contribution per hour of labor.

3. Company can provide overtime rate as long as there is no loss. So in other words labor rate per hour can be increased upto contribution per labor hour + existing rate

Max labor rate

A = $22 + $4 = $26

B= $10.5 + $4 = $14.50

C = $10.625 + $4 = $14.625

Feel free to ask any queries..

Also plz upvote it means a lot.. thank you